Let u = -71 - 9j and v = -3i +3j. Find ū+v Select the correct answer below: O 4i – 12 O 4i +12j O 10i+6j 0 -101 - 6j
Answer #1

Given a = -77 -97 and P = i +3] We have to find t he ů tv = (-77-99)+(-37 +33) = (-7+(-3))i + (-9+3); 67-3dē +(-6) = -loi-oh